Isabella invested $500 at 6% annual
interest, compounded quarterly. The
value, A, of an investment can be
calculated using the equation
A=P(1+r/n)^nt where P is the initial
investment, r is the interest rate, n is
the number of times the interest is
compounded each year, and t is time
in years. Exactly how long
will it take for her investment to be
worth four times as much (quadruple)in value?

Answer by stanbon(75887)

Isabella invested $500 at 6% annual
interest, compounded quarterly. The
value, A, of an investment can be
calculated using the equation
A=P(1+r/n)^nt where P is the initial
investment, r is the interest rate, n is
the number of times the interest is
compounded each year, and t is time
in years. Exactly how long
will it take for her investment to be
worth four times as much (quadruple)in value?
--------------------------------
Solve for "t":
4P = P(1+(0.06/4))^(4t)
(1+(0.06/4))^(4t) = 4
(1.015)^(4t) = 4
---
4t = ln(4)/ln(1.015)
t = 23.28 years
==================================
Cheers,
Stan H.
=================